OpenStudy (sambam):

Choose the correct conic section to fit the equation? (x - 3)^2 + (y - 2)^2 = 16 is it Circle Ellipse Parabola Hyperbola

OpenStudy (luigi0210):

Come on Miss Sam you know this :P

OpenStudy (sambam):

Okay how do i figure this out?

OpenStudy (luigi0210):

well they have the same values

OpenStudy (luigi0210):

it can't be a parabola b/c there is two squared values

OpenStudy (luigi0210):

and not a hyperbola b/c it's +

OpenStudy (luigi0210):

so it could either be a circle or an ellipse

OpenStudy (sambam):

Is it going to be a ellipse?

OpenStudy (luigi0210):

hm, no, if it was an ellipse it would have different values

OpenStudy (zzr0ck3r):

what would you say a^2+b^2 = 4^2 is a equation for?

OpenStudy (sambam):

i dont know @zzr0ck3r

OpenStudy (zzr0ck3r):

well you should know the answer now:)

OpenStudy (sambam):

okay well @Luigi0210 that means circle is the answer.

OpenStudy (luigi0210):

Ding ding!

OpenStudy (dan815):

ciirrclleeeeee

OpenStudy (dan815):

with radius 4 and center 3,2

OpenStudy (dan815):



OpenStudy (dan815):

iam sitting so crooked those lines look straight to me

OpenStudy (luigi0210):

well I hope we helped!

OpenStudy (sambam):

Yes little by little I think im getting it

OpenStudy (luigi0210):

good! call if you need any more help :P

OpenStudy (sambam):

\[\frac{ x^{2} }{ 9} + \frac{ y^{2} }{4}=1\] Would this be a ellipse? @Luigi0210

OpenStudy (luigi0210):

why yes, yes it would be

OpenStudy (sambam):

Okay thats great news already.

OpenStudy (dan815):

lol luigi u joker

OpenStudy (dan815):

"ive got the mind of a 13 yr old"

OpenStudy (dan815):

@Luigi0210

OpenStudy (luigi0210):

ha hey that part is true :P

OpenStudy (dan815):

haha, i have the mind of a 10 year old

OpenStudy (dan815):

so i mean the maturity of a 10 yr old